OpenStudy (anonymous):

Can someone help me solve for x: 8^(x+2) = 32, please? Thank you.

OpenStudy (gw2011):

I hope I can help. I got to the answer of x by using natural logs (ln). Specifically, the following is how I arrived at the answer: 8^(x+2)=32 (8^x)(8^2)=32 This simplifies the equation ln[(8^x)(8^2)]=ln32 ln8^x+ln8^2=ln32 xln8+2ln8=ln32 The above follow the rules of natural logs x(2.0794)+2(2.0794)=3.4657 These values came from a log table 2.0794x+4.1588=3.4657 2.0794x=3.4657-4.1588 2.0794x=-0.6931 x=-0.6931/2.0794 x=-0.333317=-0.3333 To verify this answer you can substitute this value into the original equation and you get the following: 8^(-0.3333+2)=8^1.6667=32.0022=32